1.    a) Recursively define a0 = 1, a1 = 3, a2 = 5 and an = 3an-2 + 2an-3 for n ³ 3. Calculate an for n = 3,4,5,6

(4 Marks)

 

b) Find f(2), f(3), f(4) and f(5) for the following recursive functions.

                 f(0) = 1

f(1) = 2

f(k) = (f(k -1))2 - f(k -2) + k2

(4 Marks)

 



1
Expert's answer
2022-04-04T16:07:22-0400

"a_n=3a_{n-2}+2a_{n-3}"

n=3

"a_3=3x3+2x1=11"

n=4

"a_4=3x5+2x3=20"

n=5

"a_5=3x11+2x5=43"

n=6

"a_6=3x20+2x11=82"


b."f(k)=(f(k-1))^2-f(k-2)+k^2"

"f(2)=4-1+4=7"

"f(3)=49-2+9=56"

"f(4)=3136-7+16=3145"

"f(5)=9891025-56+25=9890994"